2. Your phone slips from your pocket as you’re going up a ladder of a high building.

Your phone falls for 2.7 seconds before you heart it hit the cement below.



a. At what height did the phone fall from? (Hint: Your phone’s initial velocity is

zero.)



b. What was the velocity of your phone when it hit the cement?


1
Expert's answer
2021-10-18T10:49:32-0400

a. The height can be determined from the kinematic formula:


"h = \dfrac{gt^2}{2}"

where "t = 2.7s" and "g = 9.8m\/s^2" is the gravitational acceleration.

Thus, obtain:


"h = \dfrac{9.8\cdot 2.7^2}{2} \approx 36m"

b. The velocity can be found using the definition of acceleration:


"v = gt = 9.8\cdot 2.7\approx 26m/s"

Answer. a)36m, b)26m/s.